OBSERVATOIRE DE GENÈVE

CH-1290 Sauverny Suisse

Tel: +41 (22) 755 26 11 Fax: +41 (22) 755 39 83

Sauverny, January 7, 2002

SoftIdeas P.O.Box 8K23 Prove Islands, TO 14501 USA

Y.^{/}ref. O.^{/}ref. PL/cf69

Subject: Our activities

Dear Sir,

Enclosed^{1}, you will find a description of our activities and a check of $149.95 corresponding to your invoice #439 dated 05/28/93 for:

• 1 PiCTEX manual, describing completely all the commands implemented by the macros written by M.J. Wichura. If the manual doesn’t respond to our demand, we will send it back, ans ask for complete refund.

• 1 PiCTEX software on 5.25 inch floppy disk. The disk is to be formatted in IBM PC compatible format.

• 5 PiCTEX Quick Reference Cards.

We’ll be happy if you could send us a detailled list of your products.

We are a software developing team, and our work involves:

Editors Full-screen editors for PC and compatibles.

Compilers Full range of compilers for PC and compatibles. It includes ADA, BASIC, C, FORTH, FORTRAN, ICON, LISP, MODULA-2, PASCAL and PROLOG compilers.

1In fact, the check will be sent by separate mail

Operating Systems Multitasking, real-time OS.

Utilities A lot of useful add-on software.

We can produce cross-reference tables of product compatibilities for various software and vendors.

gnuBASE IV 419.00 nul  ok
GNATUS 123 489.95 complete N.A
Turbo Foo  99.95 partial good

We can also furnish scientific text processing systems for your computer, or for your site (Education licenses!) Here are some examples of our outputs:

x = 17y  (1)
y > a + b + c + d + e + f + g + h + i + j +
k + l + m + n + o + p  (2)
x = \{ \begin{matrix}y & if y > 0 \\ z + y & otherwise \end{matrix}

et voila … that’s it … Remember we order today:

1. 1 PiCTEX manual, describing completely all the commands implemented by the macros written by M.J. Wichura.

2. 1 PiCTEX software on 5.25 inch IBM PC compatible formatted floppy disk.

3. 5 PiCTEX Quick Reference Cards.

mm

Sincerely